11. Niles Co. has the following data related to an item of inventory: Inventory, March 1 150 units @ $1.75 Purchase, March 711. Niles Co. has the following data related to an item of inventory:Inventory, March 1150 units@ $1.75Purchase, March 7600 units@ $2.00Purchase, March 16280 units@ $2.15Sale, March 31520 unitsThe value assigned to cost of goods sold if Niles uses perpetual FIFO is
Answer #1

Cost of Goods Sold under Fifo

= (150*1.75) + (520-150)*2

= (150*1.75) + (370*2)

= 1002.50 (or) 1003

​​​​​​